in a game show the contestant is shown is 10 boxes,3 of which contain prizes.If a contestant is allowed to select any three boxes then what is the probability that
(1):- The contestant wins all the prizes.
(2):- Only one selected box contains prize.
1
Expert's answer
2012-05-31T11:36:56-0400
(1):- The contestant wins all the prizes.

The probability that the contestant wins all the prizes is

P = 1/C(10,3) = 0.0083.

(2):- Only one selected box contains prize.

P = C(3,1)*C(7,2)/C(10,3) = 0.525.